# Flow of Reasoning: Training LLMs for Divergent Problem Solving with Minimal Examples

### v4

@@ C.6 Rubik’s Cube
Additional details for baselines. In addition to IO and CoT, we also compare our approach with Hypothesis Search which belongs to discrete program search methods (Barke et al., 2024; Xu et al., 2023a; Lee et al., 2024).
@@
C.8 Logical Reasoning.
OOD data creation
We separate the in-distribution and OOD data by topics and ontology. We use the animal-related problems as in-distribution examples and the number-related problems as OOD examples.
Setup.
We use LoRA to train the model with $r=8$, $\alpha=32$, dropout=0.1. We load the LLM in fp16, and set the hyperparameters as follows: batch size = 4, learning rate = 5e-6, number of epochs = 40, and the reward weight $w=100$. See Table 13 for the prompt template of the logical reasoning task.
Additional details for Baselines.
Apart from CoT, ToT, and RAP, we compare FoR with STaR (Zelikman et al., 2022), which uses online sampling to filter our positive examples consistent with ground truth trajectories to finetune the LLM. Note that this is an easier setting than FoR which doesn’t have access to ground truth solutions. It also indicates an upper bound of SFT methods that do not rely on ground truth solutions, like. All baselines use Llama3 8B as the base model.
Appendix D Exploration and Training
@@
    Online training: (1) we employ the online policy $P_{F}(a_{t}|s_{t-1},\alpha)$, and its tempered version (2) Similar to $\epsilon$-greedy, we sample action at step $t$ by $P_{F}$ with probability $\epsilon$, and sample with uniform distribution over action space $P_{U}(a_{t}|s_{t-1})$ with $(1-\epsilon)$ probability. (3) To further explore the high-reward region, we modified the local search (Kim et al., 2023; Zhang et al., 2022). More specifically, we select the trajectory with the highest reward in a batch and conduct a destroy and reconstruction process for augmenting the trajectories to enable a higher probability of sampling successful trajectories, referring to Appendix E for more details.
@@
1:Input: $I$: number of iterations, $P_{F}$: initial LLM policy, $\mathcal{D}$: Prioritized Replay Buffer. $M$: Batch-size, $\delta$: online-offline ratio, $\mathcal{E}$: Training Dataset, $\mathcal{O}$: offline Data
2:Output: Trained policy $P_{F}$.
3:for $i=1$ to $I$ do
4:     Sample from training dataset $\mathcal{E}$ with initial state $s_{0}$ and goal $g$
5:     Sample from $u\sim[0,1]$
6:     if $u<\delta$ then
7:         // Exploration
8:         Sample $M$ online trajectories $\{\tau_{1},...,\tau_{M}\}$ with forward policy $P_{F}$
9:         Select trajectory $\tau_{m}\in\{\tau_{1},...,\tau_{M}\}$ with the largest $R(\tau_{m})$
10:         $\{\tau_{1^{\prime}}^{\prime}...\tau_{N^{\prime}}^{\prime}\}$leftarrow Local Search($\tau_{m}$)
11:         Update $\mathcal{D}\leftarrow\mathcal{D}\cup\{\tau_{1},...,\tau_{M}\}\cup\{\tau_{1^{\prime}}^{\prime}...\tau_{N^{\prime}}^{\prime}\}$
12:     else
13:         // Exploitation
14:         if is Game24 then
15:              sample $M$ offline trajectories from Offline Data $\mathcal{O}$
16:         else
17:              sample $M$ offline trajectories from $\mathcal{D}$
18:         end if
19:     end if
20:     Exploit $M$ (with $N^{\prime}$) trajectories to compute objective function in Eq 7.
21:     Update the parameter in $P_{F}$ with respect to Eq 7
22:end for
23:return $P_{F}$
Appendix E Modified Local Search
@@
    $\tau_{destroy}=(s_{0}\to\ldots\to s_{n-K}^{\prime}),\quad\quad\tau_{recon}=(s_{n-K}^{\prime}\to\ldots\to s_{n}^{\prime})$ (16)
@@
Appendix F Case Study
@@
According to Figure  4, we use the problem (3,4,6,11) to show how FoR achieves such high performance while focusing on diversity. As illustrated in the figure, we compare trajectories sampled 20 times by both SFT and FoR. While both methods produce diverse trajectories initially, FoR demonstrates better capability in reaching successful final steps from various middle steps. For example, FoR successfully transitions from intermediate steps (3,6,15) to target 24, whereas SFT fails to do so. This highlights the effectiveness of FoR’s design in simultaneously promoting diversity and ensuring accuracy.
Better robustness due to exploratory nature.
According to Figure  5, we use the problem (7,9,9,13) to demonstrate the robustness of FoR. As shown in the figure, SFT repeatedly fails by getting stuck in a single second state of (9,3,16) 20 times, while FoR successfully discovers multiple diverse trajectories leading to the correct solution. This robustness can be attributed to the exploratory nature of FoR’s training objective, which encourages the model to sample diverse successful trajectories. By expanding the search space through high-reward exploration, FoR increases the chance of finding successful outcomes. This capability not only improves the robustness of the model but also enhances its generalization to new scenarios, showcasing the effectiveness of FoR in addressing complex reasoning tasks.
Other examples
Figure 6 shows generated samples for the BlocksWorld, Figure 7 for Game24, and Table 12 for PrOntoQA, respectively.
Figure 6: Example of BlocksWorld for 6-step planning.
Figure 7: Example of game of 24.
Table 12: Examples for PrOntoQA.
Query: True or false: 31 is not imaginary. (OOD)
State Action
31 is a natural number. Natural numbers are integers.
31 is an integer. Integers are real numbers.
31 is a real number. Real numbers are not imaginary.
31 is not imaginary. Finish.
Query: True or false: Wren is not bony. (In-distribution)
State Action
Wren is a painted lady. Each painted lady is a butterfly.
Wren is a butterfly. Each butterfly is a lepidopteran.
Wren is a lepidopteran. Each lepidopteran is an insect.
Wren is an insect. Each insect is an arthropod.
Wren is an arthropod. Each arthropod is not bony.
Wren is not bony. Finish.
Figure 8: Example of GSM8K.
@@
I am playing with a set of blocks where I need to arrange the blocks into stacks. Here are the actions I can do
@@ 2. Colors in the Cube are represented in numbers: [0, 1, 2, 3, 4, 5]
Prompt for 1D-ARC
[STATEMENT]
You have the following input-output pairs:
[Initial Grid State]:
<init_state>
<python_function>
Now, please choose one function from the above list:
<action>
Figure 12: Prompt template for abstraction reasoning task.
@@
Appendix G Limitations and Future Work
Due to resource constraints, our experiments use language models with up to 13B parameters. However, we expect FoR to hold for larger models as well, and may potentially benefit larger models even more. Recent works (Xi et al., 2024; AlKhamissi et al., 2023) that finetune larger models to improve their reasoning ability with maximizing objectives usually need a large amount of data, and our data-efficient FoR may improve this process. Future work should further address two limitations in FoR.
The first is aquisation of a large amount of trajectories efficiently. Online sampling with LLMs is computationally expensive, leading to more efficient and effective strategies for exploring more complex settings such as real-world settings AlfWorld (Shridhar et al., 2020) and TravelPlanner (Xie et al., 2024a) to be further studied.
The second is faciliating FoR long-range steps reasoning. LLMs fall short in long-range planning and reasoning, thus methods like MCTS (Feng et al., 2023) or automatic reasoning system (Trinh et al., 2024) can be combined with LLMs for long-horizon divergent reasoning.
Appendix H Broader Impact
This study introduces FoR, a methodology that trains LLMs as policies to solve complex reasoning problems with better creativity and diversity. We believe this work connects LLM reasoning with GFlowNets and contributes to the application of GFlowNets to LLMs.

### v6

@@ C.6 Rubik’s Cube
Additional details for baselines. In addition to IO and CoT, we also compare our approach with Hypothesis Search which belongs to discrete program search methods (Barke et al., 2024; Xu et al., 2023a; Lee et al., 2024a).
@@
OOD data creation. We separate the in-distribution and OOD data by topics and ontology. We use the animal-related problems as in-distribution examples and the number-related problems as OOD examples.
C.8 Logical Reasoning
Setup. We use LoRA to train the model with $r=8$, $\alpha=32$, dropout=0.1. We load the LLM in fp16, and set the hyperparameters as follows: batch size = 4, learning rate = 5e-6, number of epochs = 40, and the reward weight $w=100$. See Table 10 for the prompt template of the logical reasoning task.
Additional details for Baselines. Apart from CoT, ToT, and RAP, we compare FoR with STaR (Zelikman et al., 2022), which uses online sampling to filter our positive examples consistent with ground truth trajectories to finetune the LLM. Note that this is an easier setting than FoR, which doesn’t have access to ground truth solutions. It also indicates an upper bound of SFT methods that do not rely on ground truth solutions, like. All baselines use Llama3 8B as the base model.
C.9 Mathematical Reasoning
Automatic Diversity Metric. To distinguish different solutions, we adapt a paraphrase-based method (Michail et al., 2024) to automatically evaluate the diversity with GPT-4o to measure the number of different solutions for each problem. The prompt is shown in Figure 13 Additional results. To further investigate the relationship between accuracy and the number of samples, we set the sampling counts to 1, 5, 10, and 20, and compared the performance under these configurations. As shown in Figure 4, FoR consistently outperforms SFT by a significant margin, ranging from 3.5% to 6.6%. The higher accuracy demonstrates that aligning with the reward distribution of solutions enables FoR to better explore the reasoning space, increasing the possibility of reaching the correct solution.
Figure 4: Accuracy of FoR and SFT across the different number of sampled solutions.
Appendix D Additional Analysis Results
Table 14: Performance with larger models and different families of models.
Method 4-step 6-step
Acc. (%) Diversity Acc. (%) Diversity
CoT 5-shot (Llama3-8B) 28.57 1.05 15.82 1.05
CoT 5-shot (Llama3-70B) 45.23 1.05 46.46 1.11
FoR (Llama3-8B) 98.41 1.27 78.44 1.33
FoR (Llama3-70B) 100.00 1.38 87.65 1.40
FoR (Qwen2.5-7B) 100.00 1.24 86.86 1.36
FoR (Qwen2.5-72B) 100.00 1.41 90.13 1.46
FoR (InternLM2.5-7B-Chat) 100.00 1.26 83.83 1.31
Performance with various LLMs.
We conduct experiments to evaluate FoR’s scalability and diversity gains using larger Llama-3-70B model and different family models, Qwen2.5 (7B & 72B) (Yang et al., 2024) and Intern2.5-7B-Chat (Cai et al., 2024) on the BlocksWorld task. Results shown in Table 14 indicate that FoR consistently improves accuracy and diversity with larger models. Compared to CoT baselines, FoR exhibits greater diversity gains as model size increases. This suggests that the gains of FoR become more pronounced as model capacity increases. Even with minimal data (15 examples), FoR yields clear improvements, highlighting its potential for robust performance across different base models.
Appendix E Exploration and Training
@@
    Online training: (1) we employ the online policy $P_{F}(a_{t}|s_{t-1},\alpha)$, and its tempered version (2) Similar to $\epsilon$-greedy, we sample action at step $t$ by $P_{F}$ with probability $\epsilon$, and sample with uniform distribution over action space $P_{U}(a_{t}|s_{t-1})$ with $(1-\epsilon)$ probability. (3) To further explore the high-reward region, we modified the local search (Kim et al., 2023; Zhang et al., 2022). More specifically, we select the trajectory with the highest reward in a batch and conduct a destroy and reconstruction process for augmenting the trajectories to enable a higher probability of sampling successful trajectories, referring to Appendix F for more details.
@@
1:Input: $I$: number of iterations, $P_{F}$: initial LLM policy, $\mathcal{D}$: Prioritized Replay Buffer, $M$: Batch-size, $\delta$: online-offline ratio, $\mathcal{E}$: Training Dataset, $\mathcal{O}$: offline Data
[…]
3:Sample from training dataset $\mathcal{E}$ with initial state $s_{0}$ and goal $g$
[…]
5:// Exploration
6:Sample $M$ online trajectories $\{\tau_{1},\ldots,\tau_{M}\}$ with forward policy $P_{F}$
7:Select trajectory $\tau_{m}\in\{\tau_{1},\ldots,\tau_{M}\}$ with the largest $R(\tau_{m})$
8:$\{\tau_{1^{\prime}}^{\prime},\ldots,\tau_{N^{\prime}}^{\prime}\}\leftarrow\text{Local Search}(\tau_{m})$
9:Update $\mathcal{D}\leftarrow\mathcal{D}\cup\{\tau_{1},\ldots,\tau_{M}\}\cup\{\tau_{1^{\prime}}^{\prime},\ldots,\tau_{N^{\prime}}^{\prime}\}$Else
[…]
11:Sample $M$ offline trajectories from Offline Data $\mathcal{O}$Else
[…]
13:Exploit $M$ (with $N^{\prime}$) trajectories to compute the objective function in Eq. 7
[…]
15:return $P_{F}$
Appendix F Modified Local Search
@@
    $\tau_{destroy}=(s_{0}\to\ldots\to s_{n-K}^{\prime}),$ (18)
    $\tau_{recon}=(s_{n-K}^{\prime}\to\ldots\to s_{n}^{\prime})$
@@
Appendix G Case Study
@@
According to Figure  5, we use the problem (3,4,6,11) to show how FoR achieves such high performance while focusing on diversity. As illustrated in the figure, we compare trajectories sampled 20 times by both SFT and FoR. While both methods produce diverse trajectories initially, FoR demonstrates better capability in reaching successful final steps from various middle steps. For example, FoR successfully transitions from intermediate steps (3,6,15) to target 24, whereas SFT fails to do so. This highlights the effectiveness of FoR ’s design in simultaneously promoting diversity and ensuring accuracy.
Better robustness due to exploratory nature. According to Figure  6, we use the problem (7,9,9,13) to demonstrate the robustness of FoR. As shown in the figure, SFT repeatedly fails by getting stuck in a single second state of (9,3,16) 20 times, while FoR successfully discovers multiple diverse trajectories leading to the correct solution. This robustness can be attributed to the exploratory nature of FoR’s training objective, which encourages the model to sample diverse successful trajectories. By expanding the search space through high-reward exploration, FoR increases the chance of finding successful outcomes. This capability not only improves the robustness of the model but also enhances its generalization to new scenarios, showcasing the effectiveness of FoR in addressing complex reasoning tasks.
Other examples Figure 14 shows generated samples for the BlocksWorld, Figure 15 for Game24, and Table 15 for PrOntoQA, respectively.
Appendix H Prompts.
Below, we detail all the prompts we used in all settings from Figure  789101112.
@@
I am playing with a set of blocks where I need to arrange the blocks into stacks.
Here are the actions I can do
@@ 2. Colors in the Cube are represented in numbers: [0, 1, 2, 3, 4, 5]
Prompt for 1D-ARC
[STATEMENT]
You have the following input-output pairs:
[Initial Grid State]:
<init_state>
<python_function>
Now, please choose one function from the above list:
<action>
Figure 11: Prompt template for abstraction reasoning task.
@@
Prompt for counting different solutions
Analyze the given problem-solving approaches and identify the number of distinct methods used to arrive at the solution. Consider variations in reasoning, calculation steps, and logical structuring. Directly return the total count of unique solution approaches without additional explanations.
Solution 1: "Question 3: Claire makes a 3 egg omelet every morning for breakfast. How many dozens of eggs will she eat in 4 weeks?
Question 3.1: How many eggs does she eat per day?
Answer 3.1: She makes a 3 egg omelet every morning for breakfast. So she eats 3 eggs per day. The answer is 3.
Question 3.2: How many days are in a week?
Answer 3.2: There are 7 days in a week. The answer is 7."
Solution 2: "Question 3: Claire makes a 3 egg omelet every morning for breakfast. How many dozens of eggs will she eat in 4 weeks?
Question 3.1: How many eggs does she eat in a day?
Answer 3.1: She eats 3 eggs in a day. The answer is 3.
Question 3.2: How many days are there in 4 weeks?
Answer 3.2: There are 4 * 7 = 28 days in 4 weeks. The answer is 28.
Question 3.3: How many eggs does she eat in 4 weeks?
Answer 3.3: She eats 3 eggs in a day. So she eats 3 * 28 = 84 eggs in 4 weeks. The answer is 84.
Question 3.4: Now we can answer the question: How many dozens of eggs will she eat in 4 weeks?
Answer 3.4: She eats 84 eggs in 4 weeks. 84 eggs is 84 / 12 = 7 dozens. The answer is 7."
Number: 2
Solution 1: Three of the 16 eggs go to Janet breakfast. Four of the remaining 13 eggs go to her muffins. That leaves 9 eggs to sell. The eggs sell for $2each,soshemakes9x2=$18. The answer is 18.
Solution 2: Janet lays 16 eggs a day. She eats three eggs for breakfast and bakes muffins with four this, so she has 16 - 3 - 4 = 9 eggs left. She sells these eggs at 2 each, so she make 9 x 2 = 18 at the market. The answer is 18.
Number: 1
Solution 1: <solution1>
Solution 2: <solution2>
Number:
Figure 13: Prompt template measures the number of different solutions.
Figure 14: Example of BlocksWorld for 6-step planning.
Figure 15: Example of game of 24.
Table 15: Examples for PrOntoQA.
Query: True or false: 31 is not imaginary. (OOD)
State Action
31 is a natural number. Natural numbers are integers.
31 is an integer. Integers are real numbers.
31 is a real number. Real numbers are not imaginary.
31 is not imaginary. Finish.
Query: True or false: Wren is not bony. (In-distribution)
State Action
Wren is a painted lady. Each painted lady is a butterfly.
Wren is a butterfly. Each butterfly is a lepidopteran.
Wren is a lepidopteran. Each lepidopteran is an insect.
Wren is an insect. Each insect is an arthropod.
Wren is an arthropod. Each arthropod is not bony.
Wren is not bony. Finish.
Figure 16: Example of GSM8K.
Appendix I Limitations and Future Work
Future work should further address the two limitations of FoR.
The first is acquisition of a large number of trajectories efficiently. Online sampling with LLMs is computationally expensive, leading to more efficient and effective strategies to explore more complex settings, such as real-world settings, AlfWorld (Shridhar et al., 2020), and TravelPlanner (Xie et al., 2024a) to be further studied.
The second is faciliating FoR long-range steps reasoning. LLMs fall short in long-range planning and reasoning, thus methods like MCTS (Feng et al., 2023) or an automatic reasoning system (Trinh et al., 2024) can be combined with LLMs for long-horizon divergent reasoning.
Future work can extend FoR to broader reasoning tasks, including reasoning on structured data (Min et al., 2024a, b), social reasoning (Jin et al., 2024; Yu et al., 2025), and multimodal reasoning (Zhang et al., 2023e; Kang et al., 2025).